### v1

@@ A.1.8 Fixed Effects Correction via CFFE
Note that CFFE is not strictly necessary for the EU15 sample. All EU15 eurozone members adopted the euro in 1999–2001, before any major economic crises. The treatment timing is balanced across the sample period, and the naive causal forest produces reasonable estimates. The EU28 sample, by contrast, includes countries that adopted during the 2008–2012 crisis period, creating confounding that CFFE corrects.
Table A.6 shows the results. The naive EU28 estimate of 4.8% is biased downward due to crisis-era adopters. After CFFE correction, the EU28 estimate recovers to 13.9% [12.5%, 15.3%] — close to the EU15 CFFE baseline of 14.7%.
@@
Method ATE 95% CI
Naive CF (EconML) 4.8%
CFFE (node-level FE) 13.9% [12.5%, 15.3%]
Notes: CFFE = Causal Forests with Fixed Effects, using node-level two-way FE residualization. This is theoretically superior to naive CF which applies no FE adjustment. Standard errors are cluster-robust at the pair level.
This convergence is reassuring. It suggests that the true euro effect is similar across EU15 and EU28 countries, and the apparent decline in the naive EU28 estimate was indeed driven by confounding from crisis-era adoption timing rather than genuine differences in euro effects. After controlling for fixed effects, the euro’s trade-creating effect is approximately 12–15% regardless of sample composition. These results further support our interpretation that average estimates are highly sensitive to the composition of treated units, reinforcing the importance of modeling heterogeneous treatment effects.
@@ A.1.9 Heterogeneity in the EU28 Sample
Figure A.6 shows the distribution of effects by country. The pattern differs from EU15, with newer eurozone members showing the largest effects: Malta (+52.4%), Cyprus (+34.4%), and Estonia (+29.4%) lead the rankings, while original eurozone members Portugal (+10.8%) and Greece (+10.9%) show the smallest effects. The overall ATE is +13.9%. Countries that adopted during the crisis period show smaller gains, aligning with economic intuition: countries that adopted during adverse macroeconomic conditions and had weaker pre-existing trade ties with the eurozone core experienced smaller benefits. Table A.7 provides the full country-level breakdown.
@@
Malta +43.3 18
Cyprus +33.3 18
Estonia +28.0 18
Luxembourg +23.8 18
Latvia +23.4 18
Slovenia +22.1 18
Lithuania +18.7 18
Germany +15.7 18
France +14.8 18
Italy +14.0 18
Netherlands +13.7 18
Slovakia +13.5 18
Spain +13.1 18
Belgium +12.5 18
Ireland +12.4 18
Austria +12.2 18
Finland +12.0 18
Greece +11.0 18
Portugal +10.9 18
Overall ATE +13.9
@@ A.1.10 Internet Adoption as a Potential Confounder
Luxembourg 0.476 29.1 73.0
Belgium 0.300 30.2 64.4
Germany 0.286 32.9 69.2
France 0.258 20.0 57.8
Italy 0.214 23.2 42.7
Spain 0.210 14.8 52.5
Netherlands 0.200 48.5 79.2
Ireland 0.194 19.4 56.0
Austria 0.185 33.1 64.0
Finland 0.143 43.8 76.0
Greece 0.124 10.4 36.7
Portugal 0.124 17.1 42.9
@@
     Internet 1999–2002 (early adoption): $r=0.15$
     Internet 2000–2015 (full post-treatment): $r=0.45$
@@
Note: This table tests whether the estimated euro effects could be confounded by internet adoption. The early-period correlation (1999–2002) tests whether countries with higher internet penetration at the time of euro adoption showed larger effects; this correlation is weak ($r=0.15$). The full post-treatment correlation is stronger ($r=0.45$), but this likely reflects reverse causality: countries that benefited more from the euro also developed faster economically, leading to higher internet adoption. Crucially, Luxembourg has the highest CATE but not the highest internet penetration in either period, while Finland and Netherlands have high internet but only average CATEs.
The appropriate test for confounding is whether countries with higher internet penetration at the time of euro adoption showed larger effects. Among eurozone members, the correlation between CATEs and internet penetration during the early adoption period (1999–2002) is weak ($r=0.15$). When we extend to the full post-treatment period (2000–2015), the correlation increases to $r=0.45$. However, this stronger correlation likely reflects reverse causality rather than confounding: countries that benefited more from the euro also experienced faster economic development, which in turn led to higher internet adoption. The early-period correlation is the appropriate test because it measures internet penetration before the euro’s trade effects could have influenced subsequent development trajectories.
@@
Crucially, the pattern of heterogeneity is inconsistent with internet driving the results regardless of which period we examine. Luxembourg has the highest CATE (0.48) but only moderate internet penetration—29% in 1999–2002 and 73% in 2000–2015, compared to the Netherlands at 49% and 79% respectively. Finland has high internet penetration (44% early, 76% full period) but one of the lowest CATEs (0.14). If internet were driving the results, we would expect these rankings to align; they do not.
@@
Figure A.8 visualizes these correlations. Panel A shows the weak early-period correlation ($r=0.15$), while Panel B shows the stronger full post-treatment correlation ($r=0.45$). The contrast between panels illustrates why the timing of measurement matters: the early-period correlation tests for confounding at the time of treatment, while the later-period correlation is contaminated by reverse causality. In both panels, Luxembourg (highest CATE) appears in the middle of the internet distribution, not at the top, while Finland (high internet) appears near the bottom of the CATE distribution.
@@
Figure A.8: Euro effect (CATE) vs internet adoption. Panel A plots average CATE against internet penetration during the early adoption period (1999–2002), showing a weak correlation ($r=0.15$). Panel B plots CATE against internet penetration over the full post-treatment period (2000–2015), showing a stronger correlation ($r=0.45$) that likely reflects reverse causality. In both panels, Luxembourg has the highest CATE but not the highest internet penetration, while Finland has high internet but low CATE.
@@
This evidence suggests that while internet adoption timing superficially resembles euro adoption, the heterogeneity patterns are inconsistent with internet being the driver of our results. The weak early-period correlation and the misalignment between internet and CATE rankings (Luxembourg highest CATE but not highest internet; Finland high internet but low CATE) indicate that the euro effects we estimate reflect genuine currency union effects rather than confounding from concurrent technological change.
@@
The literature on internet and trade finds modest effects. Freund and Weinhold (2004) estimate that a 10 percentage point increase in internet growth raises export growth by 0.2 percentage points. More recent work using 1996–2014 data finds elasticities of 0.03–0.13%, with larger effects among high-income countries (Rodriguez-Crespo et al., 2021). Even at the upper bound, these magnitudes are too small to explain our estimated euro effects of 12–48%. While we find moderate correlation between internet growth and CATEs ($r=0.45$), this likely reflects selection: countries more open to technological adoption may also benefit more from monetary integration, rather than internet directly driving trade gains.
@@
As a more direct test, we re-estimate our causal forest model adding pair-level internet penetration (the average of both countries’ internet usage) as a fourth effect modifier alongside GDP, GDP per capita, and pre-euro trade intensity. If internet adoption were confounding our estimates, controlling for it should substantially reduce the estimated euro effect. Instead, the ATE changes minimally: from 0.205 (22.8%) in the original specification to 0.198 (21.9%) with internet included—a reduction of only 3.4%. The confidence intervals overlap almost entirely ([0.049, 0.362] vs [0.044, 0.352]). Moreover, the individual CATEs from both models are highly correlated ($r=0.94$), indicating that the heterogeneity patterns are robust to controlling for internet. While internet penetration does have substantial feature importance (0.48) in the augmented model—suggesting it captures meaningful variation in the data—this variation is largely orthogonal to the euro effect rather than confounding it. Table A.9 presents the full comparison, and Figure A.9 visualizes these results.
@@
We also conduct this robustness check using our CFFE estimator on the EU28 sample. Here, adding internet as a fourth covariate reduces the ATE from 0.130 (13.9%) to 0.093 (9.8%)—a more substantial 28% reduction. This larger sensitivity in the CFFE specification likely reflects the expanded EU28 sample, which includes Eastern European countries where internet adoption and EU/euro accession occurred more simultaneously. The confidence intervals still overlap ([0.118, 0.142] vs [0.081, 0.105]), and the effect remains statistically significant and economically meaningful. Importantly, even after controlling for internet, the euro effect remains positive and significant at approximately 9.8%, consistent with genuine currency union benefits. Table A.10 presents the CFFE comparison.
@@
Log GDP 0.294 0.143
Log GDP per capita 0.644 0.320
Pre-euro trade 0.062 0.055
Internet penetration 0.481
@@
CATE correlation 0.94
ATE change $3.4\%$
Notes: Original model uses GDP, GDP per capita, and pre-euro trade intensity as effect modifiers. The augmented model adds pair-level internet penetration (average of both countries’ internet usage). The high CATE correlation (0.94) and minimal ATE change ($3.4\%$) indicate that internet adoption does not confound the estimated euro effects.
@@
Figure A.9: Comparison of causal forest estimates with and without internet as a covariate. Panel A shows the ATE is virtually unchanged (22.8% vs 21.9%). Panel B shows feature importance: internet has high importance but does not displace the other covariates. Panel C shows CATEs from both models are highly correlated ($r=0.94$), confirming that controlling for internet does not alter the heterogeneity patterns.
@@ A.1.11 Trade Diversion
We test for trade diversion by estimating the euro’s effect on eurozone countries’ trade with non-eurozone partners. Using our EU15 data, we examine whether eurozone countries reduced trade with the UK, Sweden, and Denmark after adopting the euro. We also use IMF Direction of Trade Statistics (DOTS) data to examine trade with non-EU partners (United States, Japan, Switzerland, Canada, Norway).
Table A.11 presents the results. The intra-eurozone effect (+46.1%) represents trade creation—the positive effect of both countries sharing the euro. For trade with non-eurozone EU members, we find a positive but imprecisely estimated effect (+15.7%, 95% CI: $59.6\%$ to $231.8\%$). For trade with non-EU partners, the effect is also positive but imprecise (+126.0%, 95% CI: $48.3\%$ to $887.5\%$).
@@
Intra-Eurozone +46.1% [+9.4%, +95.1%] 1,159 55
Eurozone → Non-Eurozone EU +15.7% [-59.6%, +231.8%] 516 24
Eurozone → Non-EU +126.0% [-48.3%, +887.8%] 1,485 55
Notes: Intra-Eurozone measures the effect of both countries adopting the euro. Eurozone → Non-Eurozone EU measures the effect of the reporter country adopting the euro on trade with non-eurozone EU members (UK, Sweden, Denmark). A negative effect would indicate trade diversion. Estimated using CausalForestDML on EU15 data, 1995-2015.
The wide confidence intervals for extra-eurozone trade reflect limited statistical power: the IMF DOTS data lacks the GDP controls available in our main Eurostat panel, and the number of non-eurozone partners is small. However, the point estimates are uniformly positive, providing no evidence of trade diversion. If anything, eurozone countries appear to have increased trade with all partners after euro adoption, though only the intra-eurozone effect is statistically significant.
This pattern is consistent with the euro reducing transaction costs for all trade involving eurozone countries, not just intra-eurozone trade. A German firm that adopts euro-denominated invoicing for French customers may also find it easier to quote prices to British or American customers in a stable currency. The absence of trade diversion suggests the euro’s trade effects are primarily trade-creating rather than trade-diverting, though we cannot rule out small diversion effects given the imprecision of our extra-eurozone estimates.

### v2

@@ A.1.8 Fixed Effects Correction via CFFE
CFFE produces substantially lower estimates than the naive causal forest for both samples: 14.1% vs 29% for EU15. This difference reflects the removal of pair-specific heterogeneity that can inflate naive estimates when high-trade pairs (which tend to have larger effects) are overweighted. The key distinction between samples is the direction of bias in naive estimates. For EU15, where all eurozone members adopted in 1999–2001 before major crises, the naive estimate is biased upward. For EU28, which includes countries that adopted during the 2008–2012 crisis period, the naive estimate is biased downward because crisis-induced trade declines are conflated with treatment effects.
Table A.6 shows the results. The naive EU28 estimate of 4.3% has a wide confidence interval due to crisis-era adopters. After CFFE correction, the EU28 estimate is 13.4% [12.1%, 14.8%] — close to the EU15 CFFE baseline of 14.1%.
@@
Method ATE Effect (%) 95% CI
Naive CF (EconML) 0.042 4.3% [-52.1%, 127.3%]
CFFE (node-level FE) 0.126 13.4% [12.1%, 14.8%]
Notes: Naive CF = Causal Forest (EconML) results from Table A.5. CFFE = Causal Forests with Fixed Effects, using node-level two-way FE residualization. ATE is the average treatment effect in log points. Effect (%) is $(\exp(\text{ATE})-1)\times 100$. Standard errors for CFFE are cluster-robust at the pair level. Sample: EU28 countries, 1995–2019.
This convergence is reassuring. It suggests that the true euro effect is similar across EU15 and EU28 countries, and the apparent decline in the naive EU28 estimate was indeed driven by confounding from crisis-era adoption timing rather than genuine differences in euro effects. After controlling for fixed effects, the euro’s trade-creating effect is approximately 13–14% regardless of sample composition. These results further support our interpretation that average estimates are highly sensitive to the composition of treated units, reinforcing the importance of modeling heterogeneous treatment effects.
@@ A.1.9 Heterogeneity in the EU28 Sample
Figure A.6 shows the distribution of effects by country. The pattern differs from EU15, with newer eurozone members showing the largest effects: Malta (+41.2%), Cyprus (+32.0%), and Estonia (+26.4%) lead the rankings, while original eurozone members Portugal (+10.7%) and Greece (+10.8%) show the smallest effects. The overall ATE is +13.4%. Countries that adopted during the crisis period show smaller gains, aligning with economic intuition: countries that adopted during adverse macroeconomic conditions and had weaker pre-existing trade ties with the eurozone core experienced smaller benefits. Table A.7 provides the full country-level breakdown.
@@
Malta +41.2 18
Cyprus +32.0 18
Estonia +26.4 18
Luxembourg +23.6 18
Latvia +22.0 18
Slovenia +21.0 18
Lithuania +17.6 18
Germany +15.3 18
France +14.5 18
Italy +13.8 18
Netherlands +13.4 18
Spain +12.9 18
Slovakia +12.8 18
Belgium +12.3 18
Ireland +12.2 18
Austria +12.0 18
Finland +11.9 18
Greece +10.8 18
Portugal +10.7 18
Overall ATE +13.4
@@ A.1.10 Internet Adoption as a Potential Confounder
Spain 0.181 14.8 52.5
Portugal 0.165 17.1 42.9
Italy 0.164 23.2 42.7
France 0.163 20.0 57.8
Germany 0.158 32.9 69.2
Greece 0.151 10.4 36.7
Netherlands 0.142 48.5 79.2
Belgium 0.133 30.2 64.4
Finland 0.121 43.8 76.0
Austria 0.116 33.1 64.0
Ireland 0.111 19.4 56.0
Luxembourg 0.099 29.1 73.0
@@
     Internet 1999–2002 (early adoption): r=0.42r=-0.42
     Internet 2000–2015 (full post-treatment): r=0.53r=-0.53
@@
Note: This table tests whether the estimated euro effects could be confounded by internet adoption. The early-period correlation (1999–2002) tests whether countries with higher internet penetration at the time of euro adoption showed larger effects; this correlation is weak. The full post-treatment correlation is stronger, but this likely reflects reverse causality: countries that benefited more from the euro also developed faster economically, leading to higher internet adoption. Crucially, Luxembourg has the highest CATE but not the highest internet penetration in either period, while Finland and Netherlands have high internet but only average CATEs.
The appropriate test for confounding is whether countries with higher internet penetration at the time of euro adoption showed larger effects. Among eurozone members, the correlation between CATEs and internet penetration during the early adoption period (1999–2002) is actually negative (r=0.42r=-0.42). The full post-treatment period (2000–2015) shows a similar negative correlation (r=0.53r=-0.53). These negative correlations indicate that countries with higher internet penetration tended to have smaller euro effects, the opposite of what we would expect if internet were driving the results.
@@
The pattern of heterogeneity is inconsistent with internet driving the results. Spain has the highest CATE (0.18) but only moderate internet penetration—15% in 1999–2002 and 53% in 2000–2015. In contrast, Netherlands has the highest early internet penetration (48%) but only an average CATE (0.14). If internet were driving the results, we would expect these rankings to align; they do not.
@@
Figure A.8 visualizes these correlations. Panel A shows the negative early-period correlation (r=0.42r=-0.42), while Panel B shows the negative full post-treatment correlation (r=0.53r=-0.53). In both panels, countries with high internet penetration (Netherlands, Finland) appear near the bottom of the CATE distribution, while countries with lower internet penetration (Spain, Portugal) show higher CATEs.
@@
Figure A.8: Euro effect (CATE) vs internet adoption. Panel A plots average CATE against internet penetration during the early adoption period (1999–2002), showing a negative correlation (r=0.42r=-0.42). Panel B plots CATE against internet penetration over the full post-treatment period (2000–2015), showing a similar negative correlation (r=0.53r=-0.53). Countries with high internet penetration (Netherlands, Finland) have lower CATEs, while countries with lower internet (Spain, Portugal) have higher CATEs.
@@
This evidence suggests that internet adoption is not confounding our results. The negative correlations and the misalignment between internet and CATE rankings indicate that the euro effects we estimate reflect genuine currency union effects rather than confounding from concurrent technological change.
@@
The literature on internet and trade finds modest effects. Freund and Weinhold (2004) estimate that a 10 percentage point increase in internet growth raises export growth by 0.2 percentage points. More recent work using 1996–2014 data finds elasticities of 0.03–0.13%, with larger effects among high-income countries (Rodriguez-Crespo et al., 2021). Even at the upper bound, these magnitudes are too small to explain our estimated euro effects of 12–48%. The negative correlation between internet and CATEs (r=0.53r=-0.53) further suggests that internet is not driving our results.
@@
As a more direct test, we re-estimate our causal forest model adding pair-level internet penetration (the average of both countries’ internet usage) as a fourth effect modifier alongside GDP, GDP per capita, and pre-euro trade intensity. If internet adoption were confounding our estimates, controlling for it should substantially reduce the estimated euro effect. Instead, the ATE changes minimally: from 0.252 (28.6%) in the original specification to 0.255 (29.0%) with internet included—a change of only 1.2%. The confidence intervals overlap almost entirely ([0.103, 0.401] vs [0.101, 0.409]). Moreover, the individual CATEs from both models are highly correlated ($r=0.94$), indicating that the heterogeneity patterns are robust to controlling for internet. Table A.9 presents the full comparison, and Figure A.9 visualizes these results.
@@
We also conduct this robustness check using our CFFE estimator on the EU28 sample. Here, adding internet as a fourth covariate reduces the ATE from 0.126 (13.4%) to 0.100 (10.5%)—a more substantial 20% reduction. This larger sensitivity in the CFFE specification likely reflects the expanded EU28 sample, which includes Eastern European countries where internet adoption and EU/euro accession occurred more simultaneously. The confidence intervals still overlap ([0.114, 0.138] vs [0.088, 0.112]), and the effect remains statistically significant and economically meaningful. Importantly, even after controlling for internet, the euro effect remains positive and significant at approximately 10.5%, consistent with genuine currency union benefits. Table A.10 presents the CFFE comparison.
@@
Log GDP 0.221 0.191
Log GDP per capita 0.050 0.027
Pre-euro trade 0.729 0.713
Internet penetration 0.069
@@
CATE correlation 0.96
ATE change $1.2\%$
Notes: Original model uses GDP, GDP per capita, and pre-euro trade intensity as effect modifiers. The augmented model adds pair-level internet penetration (average of both countries’ internet usage). The high CATE correlation and minimal ATE change indicate that internet adoption does not confound the estimated euro effects. Sample: EU15 countries, 1995–2015.
@@
Figure A.9: Comparison of causal forest estimates with and without internet as a covariate. Panel A shows the ATE is virtually unchanged (28.6% vs 29.0%). Panel B shows feature importance: internet has high importance but does not displace the other covariates. Panel C shows CATEs from both models are highly correlated ($r=0.94$), confirming that controlling for internet does not alter the heterogeneity patterns.
@@ A.1.11 Trade Diversion
We test for trade diversion by estimating the euro’s effect on eurozone countries’ trade with non-eurozone EU partners. Using our EU15 data, we examine whether eurozone countries reduced trade with the UK, Sweden, and Denmark after adopting the euro.
Table A.11 presents the results. The intra-eurozone effect (+35.5%) represents trade creation—the positive effect of both countries sharing the euro. For trade from eurozone to non-eurozone EU members, we find a positive but imprecisely estimated effect (+5.5%, 95% CI: -63.1% to +201.7%). For trade from non-eurozone EU members to eurozone countries, the effect is also positive (+34.9%, 95% CI: -2.2% to +86.1%).
@@
Intra-Eurozone +35.5% [-0.4%, +84.4%] 1,155 55
Eurozone → Non-EZ EU +5.5% [-63.1%, +201.7%] 504 24
Non-EZ EU → Eurozone +34.9% [-2.2%, +86.1%] 189 9
Notes: Intra-Eurozone measures the effect of both countries adopting the euro (trade creation). Eurozone → Non-EZ EU measures the effect of the reporter country adopting the euro on trade with non-eurozone EU members (UK, Sweden, Denmark). Non-EZ EU → Eurozone measures the reverse direction. A negative effect in the cross-eurozone pairs would indicate trade diversion. Estimated using CausalForestDML on EU15 data, 1995–2015.
The wide confidence intervals for cross-eurozone trade reflect limited statistical power: the number of non-eurozone EU partners (UK, Sweden, Denmark) is small. However, the point estimates are uniformly positive, providing no evidence of trade diversion. If anything, eurozone countries appear to have maintained or increased trade with non-eurozone EU partners after euro adoption, though only the intra-eurozone effect is precisely estimated.
This pattern is consistent with the euro reducing transaction costs for all trade involving eurozone countries, not just intra-eurozone trade. A German firm that adopts euro-denominated invoicing for French customers may also find it easier to quote prices to British or Swedish customers in a stable currency. The absence of trade diversion suggests the euro’s trade effects are primarily trade-creating rather than trade-diverting, though we cannot rule out small diversion effects given the imprecision of our cross-eurozone estimates.